OpenStudy (anonymous):

one side of a rectangle is 15 inches and the other side is x inches, what values of x will make the perimeter at most 38? length 15 width 4?

OpenStudy (anonymous):

p=2(l+b) 38=2(15+x) 19=15+x x=4 inch